The profit from the production and sale of specialty golf hats is given by the function P(x) 20x - 2000 where x is the number of hats produced and sold. (a) Producing and selling how many units will give a profit of $6000? (b) How many units must be produced and sold to avoid a loss? (a) Producing and sellingunits will give a profit of $6000 (b) To avoid a loss, units must be produced and sold.
Answer #1

ANSWER:

A) In order to get a profit of $6,000 profit p(x) should be equal to $6,000

6,000 = 20x - 2,000

6,000 + 2,000 = 20x

8,000 = 20x

x = 8,000 / 20 = 400

so selling 400 hats will give a profit of $6,000

B) In order to avoid a loss , p(x) should be equal to zero.

0 = 20x - 2,000

2,000 = 20x

x = 2,000 / 20 = 100

so selling 100 or more hats will help him in avoid the loss.